# EarthGPT-X: Enabling MLLMs to Flexibly and Comprehensively Understand Multi-Source Remote Sensing Imagery

### v1

@@ V-A2 Evaluation Metrics
We adopt Semantic Intersection over Union (S-IOU) and Semantic Similarity (SS)[73] as evaluation metrics on referring object classification tasks. Furthermore, we evaluate the performance of the proposed model on the region captioning tasks based on metrics for assessing language generation quality, including BLEU[74], METEOR[75], ROUGE[76], CIDER[77], and SPICE[78]. Those metrics provide a diverse and comprehensive assessment of the quality of identified object information and generated region captions.
TABLE II: Results on referring object classification of EarthGPT-X and other multiple MLLMs on DIOR-RSVG test set.
Methods Publication Year Formats SS S-IoU
@@
GeoChat [9] CVPR 2024 Coor 79.59 68.80
Sphinx [79] Arxiv 2023 Coor 93.72 89.37
EarthGPT [8] TGRS 2024 Coor 94.64 90.16
Natural Visual
Prompting MLLMs
Sphinx-V [50] Arxiv 2024 Box 89.07 81.62
Vip-LLava-7b [80] CVPR 2024 Box 72.56 55.94
Vip-LLava-13b [80] CVPR 2024 Box 74.51 60.53
RS Visual
Prompting MLLM
EarthMarker [22] Arxiv 2024 Point 95.96 93.49
EarthMarker [22] Arxiv 2024 Box 98.37 97.24
EarthGPT-X(Ours) Point 96.02 94.12
EarthGPT-X(Ours) Box 98.60 97.83
EarthGPT-X(Ours) Free-form 98.30 96.56
TABLE III: Results on region captioning of EarthGPT-X and other MLLMs on the OPT-RSVG test set.
Method Publication Year Format BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE CIDER SPICE
@@
Qwen-VL-Chat [9] Arxiv 2023 Coor 12.78 6.47 2.77 1.24 5.97 17.06 13.15 6.60
GeoChat [81] CVPR 2024 Coor 15.02 8.49 5.21 3.29 9.24 25.65 42.86 16.10
Sphinx [79] Arxiv 2023 Coor 22.79 15.27 10.97 7.84 12.95 30.28 105.85 20.18
EarthGPT [8] TGRS 2024 Coor 30.28 19.99 13.58 8.73 15.29 31.52 88.92 21.05
@@
Sphinx-V [50] Arxiv 2024 Box 23.35 14.84 10.37 7.16 14.10 32.08 81.79 22.46
Vip-LLava-7b [80] CVPR 2024 Box 14.67 7.29 4.06 2.25 8.20 20.17 23.47 12.79
Vip-LLava-13b [80] CVPR 2024 Box 16.39 8.40 4.77 2.80 8.47 21.36 25.85 12.41
GLaMM [80] CVPR 2024 Box 12.50 6.06 3.25 1.74 7.98 22.01 23.85 12.40
@@
Figure 3: The referring object classification results on multi-source RS images demonstrate the superior region-level RS visual understanding capability of EarthGPT-X, compared to other vanilla and visual prompting MLLMs (the black text indicates consistency with the ground truth, the red text represents incorrect answers).
@@
Figure 4: Example visualization of EarthGPT-X for multi-source (optical, SAR, and infrared) image comprehension and multiple visual tasks reasoning under various visual prompts.
V-B Referring Object Classification
To assess the merits of the proposed EarthGPT-X, we report its performance and compare it with the recent vanilla MLLMs and visual prompting MLLMs on DIOR-RSVG[7]. For referring object classification tasks, the text instruction is “Please identify the object category of the marked region in the image”. The vanilla MLLMs do not support drawing visual prompts on the images, therefore, the coordinates contained in the text instruction serve as the referring regions. The vanilla MLLMs including natural MLLMs Sphinx [79], RS MLLMs GeoChat [81], and EarthGPT [8]. We select the most recent natural visual prompting MLLMs including Sphinx-V [50], Vip-LLava [80], and the RS visual prompting MLLM model EarthMarker[22] as the comparison.
[…]
Figure 5: Pixel grounding examples of EarthGPT-X on multi-source imagery (EarthGPT-X’s interpretation of the object classes is also displayed on the figures).
V-C Region Captioning
For the brief region captioning, the test set of OPT-RSVG [82] is employed for zero-shot evaluation. We elaborately choose both natural and RS-specific MLLMs for performance comparison. Natural scene MLLMs contains Qwen-VL-Chat [9], Sphinx [79], Sphinx-V [50], Vip-LLava [80] and GLaMM [80]. Popular RS-specific MLLMs GeoChat [81] and EarthGPT [8] are selected. The text instruction for the region captioning task is designed as “Please provide a brief caption of each marked region in the image”. As displayed in Tab. III, the proposed EarthGPT-X significantly outperforms the previous works on all of the evaluation metrics. Notably, using boxes as visual prompts can yield the best performance on region captioning and dialogue ability.
V-D Comprehensive Capabilities Evaluation
Traditional vision models in the RS field primarily focus on a single visual modality and single tasks, exhibiting poor generalization[83, 84, 85]. Thanks to our M-RSVP’s rich multi-sensor knowledge, EarthGPT-X has robust generalization capabilities across optical, SAR, and infrared visual modalities. To verify and assess the comprehensive capabilities of the proposed model, we conduct a thorough qualitative evaluation. Our assessment also covers multi-source images and multiple levels of image interpretation granularity, including image, point, box, and free-form prompted levels. Moreover, the evaluated visual tasks are extensive, such as brief captioning, detailed captioning, ground captioning, region captioning, referring object classification, and inter-relationship analyses. As displayed in Fig. 4, our model demonstrates remarkable visual reasoning and analytical abilities. EarthGPT-X can provide detailed descriptions for multi-source images, as shown in the interpretation results of the first image, which includes overall image depiction, number of targets, target orientation, and scene reasoning. Notably, in the relationship analysis of the second and third images, EarthGPT-X provides detailed and comprehensive responses. Specifically, EarthGPT-X adeptly summarizes the connections between various regions and the overall image relationship, and then analyzes the characteristics and attributes of the targets within each specific area. In summary, EarthGPT-X demonstrates a profound ability to dissect and understand complex visual information in the multi-source RS MLLMs.
V-E Pixel Grounding
To evaluate the pixel grounding capabilities of the proposed model, we specifically chose images that feature complex and different scenes with multiple targets. This experiment aims to test the model’s robustness and precision in identifying and segmenting diverse elements within a single frame. As shown in Fig. 5, the first test image is the port scene, we mark a ship and a tiny vehicle to assess the model’s ability to discern and accurately interpret tiny targets amidst a busy background. Surprisingly, EarthGPT-X not only recognized these smaller objects but also effectively segmented them from the surrounding environment. Another test involved an image from an urban sports center where we marked five distinct targets scattered across different areas, including varied elements such as a swimming pool, tennis court, vehicle, etc. This scenario tests the model’s capacity to handle spatially separated objects within a complex urban setting. EarthGPT-X demonstrates exceptional performances in precise object recognition and pixel-grounding. Moreover, these tests highlight EarthGPT-X’s advanced pixel understanding capabilities and its potential to enhance RS image analyses across different visual modalities.
V-F Ablation Study
In this section, we analyze the impact of key components of EarthGPT-X in detail.
TABLE IV: Ablation results on different visual encoders.
CLIP CLIP DINOv2 SS S-IoU
ConvNeXt-L RN50$\times$16 ViT-L/14
@@ V-F Ablation Study
TABLE V: Ablation results on Image Pre-processing.
Settings Visual Encoder SS S-IoU
@@
Multi-scale + Sub-images MVE 98.40 97.13
@@ V-F1 Different Visual Encoders
To show the effectiveness of the proposed hybrid MVE module, we ablate on the performance of our EarthGPT-X when using different visual encoders. Specifically, this ablation uses CLIP-ConvNeXt-L, CLIP RN50×16, and DINOv2 ViT-L/14 for comparison. The results are shown in Tab. IV, it is clear that the setting using the ConvNeXt-L and DINOv2 ViT-L/14 as the mixed visual experts achieve the most superior performance. The performance of both SS and S-IoU increases accordingly. These results demonstrate that utilizing a network architecture with complementary functions has deepened the understanding of local information and contextual dependencies on RS imagery.
@@ V-F2 Sub-images Decomposition and Multi-scale
In the visual feature extraction phase of our approach, we enhance the integration of visual features by incorporating multiple scales and high-resolution sub-images, enabling a more profound comprehension of RS images. To evaluate the effectiveness of our modified feature extraction strategy, we conduct ablation studies focusing on the impact of multi-scale features and high-resolution sub-images on the performance of our EarthGPT-X model. As shown in Tab. V, utilizing the multi-scale and sub-images features setting achieves superior performance over single-scale features. These results demonstrate that the detailed analysis of these high-resolution sub-images and more scales allows for capturing finer details that are crucial for accurate interpretation and analysis.
VI Conclusion

### v2

@@ V-A2 Evaluation Metrics
We adopt three categories of evaluation metrics tailored to different visual tasks in our benchmark. (i) Region and Image Captioning tasks: We employ standard language generation metrics, including BLEU [81], METEOR [82], ROUGE [83], CIDEr [84], and SPICE [85]. (ii) Scene Classification: We report classification accuracy, which directly reflects the model’s ability to recognize high-level scene semantics in RS images. (iii) Referring Object Classification: We employ two semantic-aware metrics: Semantic Intersection over Union (S-IOU) and Semantic Similarity (SS) [86, 3, 77]. SS measures the similarity of predicted and ground-truth labels in a semantic embedding space, while S-IOU captures the semantic overlap between them by computing the intersection-over-union in the embedding space. Those metrics provide a comprehensive assessment of the quality of the identified object information and generated captions. Consider the referring expression “A airplane is on the right of a small vehicle”, where the ground-truth target labels are airplane and vehicle. Suppose the model predicts aircraft and car, both outputs achieve high SS since they are semantically close to the labels in the embedding space, and also high S-IoU because of strong token-level overlap. In contrast, if the predictions are ship and building, SS drops due to weaker semantic similarity, while S-IoU becomes very low as the key object words differ significantly from the ground-truth labels.
TABLE III: Zero-shot Scene Classification (UCMerced and AID).
Models UCMerced Acc AID Acc
Qwen-VL [87] 62.90 52.60
MiniGPTV2 [88] 4.76 12.90
LLaVa-1.5 [29] 68.00 51.00
Sphinx [89] 62.76 58.20
InternVL [90] 58.23 60.40
GeoChat [9] 84.43 72.03
EarthMarker [15] 86.52 77.97
EarthGPT-X (Ours) 87.89 78.09
TABLE IV: Referring object classification (DIOR-RSVG).
Models Point Box (Coor) Free-form
$S$ S-IoU $S$ S-IoU $S$ S-IoU
@@
InternVL2.5 [38] 54.54 37.91
InternVL3 [91] 50.84 26.99
GeoChat [9] 79.59 68.80
Sphinx [89] 93.72 89.37
EarthGPT [8] 94.64 90.16
Visual Prompting MLLMs
Sphinx -V [77] 86.21 78.78 89.07 81.62 88.46 79.97
Vip-LLava-7b [18] 68.32 51.98 72.56 55.94 71.24 53.99
Vip-LLava-13b [18] 70.02 57.61 74.51 60.53 74.05 59.83
EarthMarker [15] 95.96 93.49 98.37 97.24
EarthGPT-X (Ours) 96.02 94.12 98.60 97.83 98.30 96.56
Notes: Since vanilla MLLMs lack visual prompt support, textual coordinates are treated as the referring regions. ‘—’ means no such capability.
TABLE V: Referring Object Classification (MSAR).
Models Formats $S$ S-IoU
InternVL2.5 [38] Coor 37.05 7.69
InternVL3 [91] Coor 32.72 2.89
Sphinx-V [77] Box 34.76 2.03
EarthGPT-X (Ours) Box 99.85 97.75
TABLE VI: Supervised Image Captioning (NWPU-captions).
Models BLEU1 BLEU2 BLEU3 BLEU4 METEOR ROUGE-L CIDEr(0-5) SPICE
Expert Models
CSMLF [92] 77.0 64.9 53.2 47.1 32.0 57.8 106.5 26.5
Qu, et. al [60] 72.5 60.3 51.8 45.5 33.6 59.1 117.9 27.6
Attention (soft) [59] 73.1 60.9 52.5 46.2 33.9 59.9 113.6 28.5
Attention (hard) [59] 73.3 61.0 52.7 46.4 34.0 60.0 110.3 28.4
FC-Att [93] 73.6 61.5 53.2 46.9 33.8 60.0 123.1 28.3
SM-Att [93] 73.9 61.7 53.2 46.8 33.0 59.3 123.6 27.6
MLCA-Net [94] 74.5 62.4 54.1 47.8 33.7 60.1 126.4 28.5
EarthGPT-X (Ours) 85.2 74.2 64.0 55.9 39.5 71.5 163.7 27.3
Figure 5: Image brief and detailed captioning results on optical, infrared, and SAR modalities, demonstrating the superior scene-level visual understanding ability of EarthGPT-X.
Figure 6: The referring object classification results on SAR and infrared RS images demonstrate the superior object-level RS visual understanding capability of EarthGPT-X, compared to other vanilla and visual prompting MLLMs (The black text indicates consistency with the ground truth, the red text represents incorrect answers).
TABLE VII: Zero-shot region captioning (OPT-RSVG).
Models Format BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE CIDEr SPICE
@@
Qwen-VL [9] Coor 12.78 6.47 2.77 1.24 5.97 17.06 13.15 6.60
InternVL2.5 [38] Coor 9.87 4.18 1.69 0.52 5.71 10.55 9.22 8.28
InternVL3 [91] Coor 15.42 8.91 6.34 4.38 18.64 14.28 13.27 14.87
GeoChat [87] Coor 15.02 8.49 5.21 3.29 9.24 25.65 42.86 16.10
Sphinx [89] Coor 22.79 15.27 10.97 7.84 12.95 30.28 105.85 20.18
EarthGPT [8] Coor 30.28 19.99 13.58 8.73 15.29 31.52 88.92 21.05
@@
Sphinx-V [77] Box 23.35 14.84 10.37 7.16 14.10 32.08 81.79 22.46
Vip-LLava-7b [18] Box 14.67 7.29 4.06 2.25 8.20 20.17 23.47 12.79
Vip-LLava-13b [18] Box 16.39 8.40 4.77 2.80 8.47 21.36 25.85 12.41
@@
Note: The underline denotes the best zero-shot performance achieved under box prompts.
TABLE VIII: Region Captioning on SAR and Infrared datasets.
Models SAR-Ship (Supervised) HIT UAV (Zero-Shot)
ROUGE-1 ROUGE-L METEOR ROUGE-1 ROUGE-L METEOR
Generic MLLMs
GPT-4o[95] 7.49 7.24 7.07 10.96 9.02 8.23
InternVL2[90] 9.67 8.67 8.19 11.00 9.53 8.40
Specialized RS MLLMs
GeoChat[9] 57.15 57.15 52.20 59.85 59.85 51.31
EarthDial[23] 63.10 63.10 54.83 61.83 61.83 52.80
EarthGPT-X (Ours) 65.25 65.25 55.90 64.40 64.40 51.56
@@
Figure 7: Qualitative examples of EarthGPT-X performing multi-task reasoning and comprehending multi-source RS imagery (optical, SAR, and infrared) under various visual prompts.
Figure 8: Illustration of multiple visual prompting strategies. Specifically, Point Prompt highlights local functional roles of individual regions, Box Prompt emphasizes global thematic coherence and spatial relations, and Free-form Prompt provides concise contextual abstraction. Despite these differences, all three prompting types guide the model to establish precise region–text alignment.
TABLE IX: Ablation results about the design of the visual encoders.
CNN Model ViT Model Classification Region Captioning Ref. Obj. Classification
@@
[…]
V-B Scene Classification
We evaluate EarthGPT-X on two RS benchmarks: AID [96] and UCMerced [97]. Following GeoChat [9], we use 20% of AID for testing, while the entire UCMerced serves as a zero-shot test set. To fit the visual prompting setting, we define a full-image box $0,0,\mathrm{width},\mathrm{height}]$ and use the instruction: “Please identify the object category of each marked region in the image.” Classification accuracy is computed under zero-shot settings. As shown in Tab. III, EarthGPT-X achieves 78.09% on AID and 87.89% on UCMerced, outperforming all baselines. This further confirms that the proposed visual prompting paradigm achieves stronger task alignment than directly applying vanilla MLLMs (e.g., Sphinx and GeoChat) without prompt adaptation.
V-C Referring Object Classification
V-C1 Evaluation on Optical Data
Referring object classification aims to identify the category labels of objects referred to by visual prompts and textual instructions. To assess the merits of the proposed EarthGPT-X, we report its performance and compare it with the recent vanilla and visual prompting MLLMs on DIOR-RSVG[7]. We adapt DIOR-RSVG by converting referring expressions and their bounding boxes into classification samples, where each target region serves as input and its category label as output. The text instruction is “Please identify the object category of the marked region in the image”. The vanilla MLLMs do not support drawing visual prompts on the images, therefore, the coordinates contained in the text instruction serve as the referring regions. The vanilla MLLMs used for comparison include natural-domain models such as Sphinx [89], InternVL2.5-1B[38], InternVL3-1B[91], as well as RS-domain models including GeoChat [87] and EarthGPT [8]. For visual prompting MLLMs, we select the most recent representatives, namely the natural-domain models Sphinx-V [77] and Vip-LLava [18], together with the RS-domain model EarthMarker [15]. We conduct experiments under three levels of visual prompting (point, box, and free-form) to provide a fair and comprehensive comparison across models. Following [77, 1], for point prompts, we randomly select a point within the ground-truth object that is also close to its boundary. For box prompts, we use the ground-truth bounding box provided by DIOR-RSVG. For free-form prompts, we randomly perturb the center positions and zoom the box prompts to simulate. As shown in Tab. IV, EarthGPT-X achieves optimal results in S-IoU under box prompts. We also observe that the box-level results are better than the point and free-form prompt ones.
V-C2 Evaluation on SAR Data
We evaluate EarthGPT-X and recent InternVL series[38, 91], Sphinx-V[77] on the MSAR[69] test set (Tab. V). EarthGPT-X, trained with SAR-specific samples, achieves superior performance (SS: 99.85%, S-IoU: 97.75%), showing strong capacity in SAR-oriented understanding. In contrast, natural MLLMs InternVL2.5 and InternVL3, without SAR training, perform significantly lower. The superior results of EarthGPT-X illustrate its stronger modality adaptation and specialization capabilities, which are crucial for accurate SAR imagery understanding in the RS applications.
V-D Image Captioning
To evaluate the image captioning capabilities, we adopt the NWPU-Captions [94] dataset to benchmark EarthGPT-X against other expert models under supervised settings. During evaluation, we use the instruction “Please provide a brief caption of each marked region in the image” alongside a full-image box $0,0,\mathrm{width},\mathrm{height}]$ as the visual prompt to guide the captioning process. As reported in Tab. VI, EarthGPT-X consistently outperforms existing expert methods across all evaluation metrics. These results indicate that by treating the entire image as a unified region and leveraging multi-domain training on both natural and RS data, EarthGPT-X is capable of generating accurate and semantically rich descriptions for RS scenes.
V-E Region Captioning
V-E1 Evaluation on Optical Data
For the brief region captioning on optical data, the test set of OPT-RSVG [98] is employed for zero-shot evaluation. We elaborately choose both natural and RS-specific MLLMs for performance comparison. Natural scene MLLMs contains Qwen-VL [9], InternVL2.5-1B[38], InternVL3-1B[91], Sphinx [89], Sphinx-V [77], Vip-LLava [18] and GLaMM [18]. Popular RS-specific MLLMs GeoChat [87] and EarthGPT [8] are selected. The text instruction for the region captioning task is designed as “Please provide a brief caption of each marked region in the image”. As shown in Tab. VII, the proposed EarthGPT-X consistently outperforms previous works across all evaluation metrics. Notably, EarthGPT-X achieves significant gains over Sphinx-V, despite the latter being trained on large-scale RS datasets. This highlights the superior performance and stronger generalization ability of EarthGPT-X in multi-source RS data understanding. Notably, using boxes as visual prompts can yield the best performance on region captioning ability.
V-E2 Evaluation on SAR and Infrared Data
As displayed in Tab. VIII, for the multi-source brief captioning task, we evaluate EarthGPT-X on two representative datasets: SAR-Ship [99] for supervised evaluation and HIT UAV [100] under a zero-shot setting. These datasets enable us to validate the generalization ability of EarthGPT-X across different modalities. For comparison, we include two representative general-purpose MLLMs, GPT-4o and InternVL2 [90]. We further reference the first RS MLLM, GeoChat [9], and the latest multi-source understanding model, EarthDial [23], as strong domain-related baselines. Experimental results demonstrate that EarthGPT-X achieves superior overall performance on multi-source datasets.
V-F Ablation Study
TABLE X: Ablation results on Image Pre-processing.
Settings Visual Encoder $S$ S-IoU
@@ V-F Ablation Study
Multi-scale + Sub-images MVE 98.60 97.83
TABLE XI: Computation Analysis and Comparisons
 Models  Inference Time(s)  GPU Memory(MiB)
MLLMs
Sphinx 1.08 41853
EarthGPT 0.79 41766
GeoChat 0.74 15544
Qwen-VL 0.35 38145
Visual Prompting Models
Sphinx-V 1.95 41794
Vip-LLava-13B 0.84 27748
EarthGPT-X(Ours) 1.60 41006
@@ V-F1 Different Visual Encoders
To show the effectiveness of the proposed hybrid MVE module, we ablate on the performance of our EarthGPT-X when using different visual encoders. Specifically, this ablation uses CLIP-ConvNeXt-L, CLIP RN50×16, and DINOv2 ViT-L/14 for comparison. The results are shown in Tab. IX, it is clear that the setting using the ConvNeXt-L and DINOv2 ViT-L/14 as the mixed visual experts achieves the most superior performance on scene classification, region captioning, and referring object classification tasks. These results demonstrate that utilizing a network architecture with complementary functions has deepened the understanding of local information and contextual dependencies on RS imagery.
@@ V-F2 Sub-images Decomposition and Multi-scale
In the visual feature extraction stage, we strengthen feature integration by leveraging multi-scale features and high-resolution sub-images for richer RS image comprehension. Ablation results in Tab. X show that this design significantly outperforms single-scale settings on the DIOR-RSVG dataset, confirming that capturing finer details through multiple scales and sub-images is crucial for accurate interpretation.
V-G Computation Analysis and Comparisons
In this section, we compare the GPU memory usage (MiB) and inference time (seconds) of EarthGPT-X with other MLLMs and visual prompting models. All evaluations are conducted on the region captioning task using an NVIDIA RTX A6000. For both MLLMs and visual prompting models, GPU memory usage is primarily attributed to model weight storage. As illustrated in Tab. XI, EarthGPT-X consumes more GPU memory than Sphinx [77], GeoChat [9], EarthGPT [8], Qwen-VL [87], and Vip-LLava-13B [18], but less than Sphinx-V [77]. Similarly, its inference time is longer than all other models except Sphinx-V, due to its larger parameter scale. However, the time difference remains within one second, indicating that the performance gap is marginal and does not affect deployment in practical applications, as EarthGPT-X remains well within the capacity of commonly used GPUs. In future work, we plan to optimize EarthGPT-X’s inference efficiency to better balance speed and interpretability.
Figure 9: Examples of EarthGPT-X performing visual reasoning and pixel-level grounding on optical imagery across diverse scenes (harbor, urban stadium, and airport). Better to zoom in to view clearly.
Figure 10: Examples of EarthGPT-X performing small vehicle recognition and reasoning with pixel-level grounding on multi-source imagery (i.e., SAR, infrared, optical–daytime, and optical–nighttime). Better to zoom in to view clearly.
VI Visualization
VI-A Multi-source Image Captioning
We present visualization results of EarthGPT-X on optical, SAR, and infrared images to demonstrate its capability in understanding diverse RS scenes. Given an entire image as the region of interest, we use the visual prompt $0,0,\mathrm{width},\mathrm{height}]$ and provide instructions such as “Please provide a brief caption of each marked region in the image. ” or “Please provide a detailed caption of each marked region in the image.” As shown in Fig. 5, the model generates accurate and semantically rich captions across modalities, effectively handling varied visual characteristics and noise patterns.
VI-B Referring Object Classification
We conduct qualitative analysis to provide an intuitive performance display of EarthGPT-X. As shown in Fig. 6, the infrared image showcases a residential leisure area, and for the three marked regions, all other MLLMs respond incorrectly. Conversely, EarthGPT-X successfully recognizes the human, vehicle, and basketball court targets. Following this is a fuzzy port SAR image, where it is difficult to visually distinguish between ships and storage tanks. EarthGPT-X delivers precise interpretations for these targets, whereas other models perform poorly, except for EarthGPT. These results prove the significance of our work that empowers the MLLMs with fine-grained reasoning ability for object-level comprehension.
VI-C Multi-level Multi-task Reasoning
Traditional vision models in the RS field primarily focus on a single visual modality and single tasks, exhibiting poor generalization[101, 102, 103]. Thanks to our M-RSVP’s rich multi-sensor knowledge, EarthGPT-X has robust generalization capabilities across optical, SAR, and infrared visual modalities. To verify and assess the comprehensive capabilities of the proposed model, we conduct a thorough qualitative evaluation. Our assessment also covers multi-source images and multiple levels of image interpretation granularity, including image, point, box, and free-form prompted levels. Moreover, the evaluated visual tasks are extensive, such as brief captioning, detailed captioning, ground captioning, region captioning, referring object classification, and inter-relationship analyses. As displayed in Fig. 7, our model demonstrates remarkable visual reasoning and analytical abilities. EarthGPT-X provides detailed descriptions for multi-source images, as evidenced by the interpretation of the first image, which includes overall scene depiction, target counts, orientations, and higher-level reasoning. In the relationship analysis of the second and third examples, the model generates comprehensive and structured responses, effectively capturing both local and global semantics. Notably, even for low-quality SAR and infrared maritime images, EarthGPT-X can infer relative spatial relations, target sizes, vessel functionalities, and operational states (e.g., docking at a port). In summary, EarthGPT-X demonstrates a profound ability to dissect and understand complex visual information in the multi-source RS MLLMs.
VI-D Multiple Visual Prompts
To evaluate the adaptability of EarthGPT-X under different prompting shapes, we investigate point prompt, box prompt, and free-form prompt. These prompts vary in the way visual regions are highlighted to the model, thereby influencing the style and granularity of responses. i) Similarities: Across all prompt types, EarthGPT-X consistently grounds the responses to the marked regions, ensuring semantic coherence and correct alignment with the visual input. ii) Differences: As shown in the Fig. 8, under the point prompt, EarthGPT-X generates a caption emphasizing the spatial relation, which reflects its local context-awareness. Under the box prompt, the model highlights the absolute spatial location, showing a region-level grounding behavior. In contrast, the free-form prompt yields a more generic description, focusing on object-level attributes (Note that these examples are qualitative illustrations, not unique or deterministic results, since different prompt placements may lead to varied outputs).
VI-E Zero-shot Pixel-level Grounding and Reasoning
All pixel grounding and reasoning experiments are conducted in a zero-shot setting, and the capability is obtained without any task-specific training.
VI-E1 Different Scenes
To evaluate the robustness under complex scenes, we select RS images with multiple diverse targets. As shown in Fig. 9, in the port scene, we mark a ship and a tiny vehicle to examine the model’s ability to localize small objects within a cluttered background. EarthGPT-X not only successfully detects these targets but also separates them from the surroundings. In another urban sports center scene, five spatially separated objects (e.g., swimming pool, tennis court, and vehicle) are interpreted. EarthGPT-X demonstrates exceptional performance in precise object recognition and pixel-grounding.
VI-E2 Different Image Modalities
As shown in Fig. 10, we further evaluate on small vehicle targets across heterogeneous image modalities, including optical, infrared, and SAR. The test data is from VisDrone-DroneVehicle [104] and SAR_Vehicle_Detection_Dataset[105]. Despite the significant domain gaps among these modalities, EarthGPT-X consistently identifies and segments small and tiny vehicles, showing its strong generalization and adaptability to heterogeneous RS image types.
VII Conclusion